OpenStudy (anonymous):

Find the slope m of the tangent line to the graph of the function at the given point and determine an equation of the tangent line. f(x)=7x-3x^2 at (-3, -48)

OpenStudy (helder_edwin):

can u find the derivative of \[ \large f(x)=7x-3x^2 \]

OpenStudy (anonymous):

7-3x

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

what about the \(x^2\)

OpenStudy (helder_edwin):

\[ \large f'(x)=7-3\cdot2x=7-6x \]

OpenStudy (helder_edwin):

the slope of the tangent line is then \[ \large f'(-3)=7-6(-3)= \]